OpenStudy (anonymous):

Find the exact value of the remaining trigonometric functions of theta. tan theta = 2*6^1/2, cos theta < 0

OpenStudy (anonymous):

\[\tan \theta = 2\sqrt{6}\]

OpenStudy (fretje):

\[\theta =\tan^{-1} \left( 2\sqrt{6} \right)\] \[\theta =78.463 +n.\Pi\]

OpenStudy (anonymous):

why you selected these 2 func. ?

OpenStudy (anonymous):

I need to find the EXACT value of sin theta, cos theta, cot theta, sec theta, and csc theta. But if anyone can help me get one of them, I think I can figure out the rest.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you can not h3

OpenStudy (fretje):

one is easy: \[\cot \theta = \left( 1 \over \tan \theta \right) \]

OpenStudy (fretje):

so \[\cot \theta = \left( 1 \over 2.\sqrt{6} \right)\]

OpenStudy (anonymous):

This has become very confused.

OpenStudy (fretje):

and Theta is in the third quadrant

OpenStudy (anonymous):

Yes, I got that.

OpenStudy (fretje):

i got something: i use the t formulas for getting sin theta

OpenStudy (fretje):

\[\tan \theta = \left( 2.t \over 1-t ^{2} \right)\]

OpenStudy (fretje):

substitute \[\tan \theta = 2.\sqrt{6}\]

OpenStudy (fretje):

and solve for t, that is a quadratic equation ( http://nl.wikipedia.org/wiki/Vierkantsvergelijking)

OpenStudy (fretje):

determinant is 100, and t1 and t2 are \[t1 =\left( 3 \over \sqrt{6} \right)\ t2 = \left( 2 \over \sqrt{6} \right)\]

OpenStudy (fretje):

then i get for cos theta \[\cos \theta = \left( 1-t ^{2} \over 1+t ^{2}\right) \]

OpenStudy (fretje):

which gives me \[\cos \theta = \left( -1\over 5 \right) and \cos \theta =\left( 1 \over 5 \right)\]

OpenStudy (fretje):

there is only one possible solution: cos theta < 0 as in the question so \[\cos \theta = \left( -1\over 5 \right)\]

OpenStudy (anonymous):

Perhaps I can try from the t formulas...thanks for your help.

OpenStudy (fretje):

last solution must be correct: \[\cos \theta = \left( -1\over 5 \right)\]

OpenStudy (fretje):

similarly i get the sin theta: \[\sin \theta = \left( 2.t \over 1+t ^{2} \right)\]

OpenStudy (fretje):

with \[t1 =\left( 3 \over \sqrt{6} \right)\ t2 = \left( 2 \over \sqrt{6} \right)\]

OpenStudy (fretje):

\[\sin \theta =\left( 12 \over 5 \sqrt{6} \right)\]

OpenStudy (anonymous):

I get (4*sqrt 6)/25

OpenStudy (anonymous):

Oh, I was supposed to divide by 2 to get t, right?

OpenStudy (fretje):

if i use the two solutions of the quadratic equation t1 and t2 i get for sin theta same answer.

OpenStudy (fretje):

\[\sin \theta = \left( 2.t \over 1+t ^{2} \right)= \left( \left( 6 \over \sqrt{5} \right)\over \left( 1+\left( 9 \over 6 \right) \right) \right) = \left( 12 \over 5.\sqrt{6} \right)\]

OpenStudy (fretje):

\[\sin \theta = \left( 2.t \over 1+t ^{2} \right)= \left( \left( 6 \over \sqrt{6} \right)\over \left( 1+\left( 9 \over 6 \right) \right) \right) = \left( 12 \over 5.\sqrt{6} \right)\]

OpenStudy (anonymous):

isn't t defined as tan X = X/2?

OpenStudy (anonymous):

Rather, t would be X/2 given tan X?

OpenStudy (fretje):

for t1, and \[\sin \theta = \left( 2.t \over 1+t ^{2} \right)= \left( \left( 4 \over \sqrt{6} \right)\over \left( 1+\left( 4 \over 6 \right) \right) \right) = \left( 24 \over 10.\sqrt{6} \right)\] for t2, which is the same result

OpenStudy (fretje):

different tangent, that is the point with the t formulas, you dont need to know what t stands for.

OpenStudy (fretje):

now sec and csc are just resp. 1/cos and 1/sin

OpenStudy (fretje):

and sin theta is negative too: \[\sin \theta = -\left( 12 \over 5.\sqrt{6} \right)\]

OpenStudy (anonymous):

Thanks!